Solution Manual for Probability and Statistical Inference, 10th Edition
- Statistics
- 143 pages
- Shared May 2025
With Solution Manual for Probability and Statistical Inference, 10th Edition, you can tackle your textbook assignments with confidence, knowing you have the correct answers.
Preview 16 of 143 pages
See the first 16 pages exactly as they appear in the document.
Page 1 Read page 1 as text
S OLUTIONS M ANUAL P ROBABILITY AND S TATISTICAL I NFERENCE T ENTH E DITION Robert V. Hogg Elliot A. Tanis Dale L. Zimmerman
Page 2 Page 3 Read page 3 as text
iii Contents Preface v 1 Probability 1 1.1 Properties of Probability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Methods of Enumeration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Conditional Probability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Independent Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.5 Bayes’ Theorem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2 Discrete Distributions 7 2.1 Random Variables of the Discrete Type . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 Mathematical Expectation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.3 Special Mathematical Expectations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.4 The Binomial Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5 The Hypergeometric Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.6 The Negative Binomial Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.7 The Poisson Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3 Continuous Distributions 19 3.1 Random Variables of the Continuous Type . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.2 The Exponential, Gamma, and Chi-Square Distributions . . . . . . . . . . . . . . . . . 26 3.3 The Normal Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.4 Additional Models . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 4 Bivariate Distributions 33 4.1 Bivariate Distributions of the Discrete Type . . . . . . . . . . . . . . . . . . . . . . . . 33 4.2 The Correlation Coefficient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.3 Conditional Distributions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.4 Bivariate Distributions of the Continuous Type . . . . . . . . . . . . . . . . . . . . . . 37 4.5 The Bivariate Normal Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5 Distributions of Functions of Random Variables 45 5.1 Functions of One Random Variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.2 Transformations of Two Random Variables . . . . . . . . . . . . . . . . . . . . . . . . 47 5.3 Several Random Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 5.4 The Moment-Generating Function Technique . . . . . . . . . . . . . . . . . . . . . . . 53 5.5 Random Functions Associated with Normal Distributions . . . . . . . . . . . . . . . . 55 5.6 The Central Limit Theorem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.7 Approximations for Discrete Distributions . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.8 Chebyshev’s Inequality and Convergence in Probability . . . . . . . . . . . . . . . . . 61 5.9 Limiting Moment-Generating Functions . . . . . . . . . . . . . . . . . . . . . . . . . . 62
Page 4 Read page 4 as text
iv Contents 6 Point Estimation 63 6.1 Descriptive Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 6.2 Exploratory Data Analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3 Order Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 6.4 Maximum Likelihood Estimation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 6.5 A Simple Regression Problem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 6.6 Asymptotic Distributions of Maximum Likelihood Estimators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 6.7 Sufficient Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 6.8 Bayesian Estimation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 7 Interval Estimation 87 7.1 Confidence Intervals for Means . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 7.2 Confidence Intervals for the Difference of Two Means . . . . . . . . . . . . . . . . . . . 88 7.3 Confidence Intervals For Proportions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.4 Sample Size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 7.5 Distribution-Free Confidence Intervals for Percentiles . . . . . . . . . . . . . . . . . . . 92 7.6 More Regression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 7.7 Resampling Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 8 Tests of Statistical Hypotheses 107 8.1 Tests About One Mean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 8.2 Tests of the Equality of Two Means . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 8.3 Tests for Variances . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 8.4 Tests about Proportions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 8.5 Some Distribution-Free Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 8.6 Power of a Statistical Test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 8.7 Best Critical Regions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 8.8 Likelihood Ratio Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 9 More Tests 127 9.1 Chi-Square Goodness-of-Fit Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 9.2 Contingency Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 9.3 One-Factor Analysis of Variance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 9.4 Two-Way Analysis of Variance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 9.5 General Factorial and 2 k Factorial Designs . . . . . . . . . . . . . . . . . . . . . . . . . 135 9.6 Tests Concerning Regression and Correlation . . . . . . . . . . . . . . . . . . . . . . . 136 9.7 Statistical Quality Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137
Page 5 Read page 5 as text
Chapter 1 Probability 1 Chapter 1 Probability 1.1 Properties of Probability 1.1-2 Sketch a figure and fill in the probabilities of each of the disjoint sets. Let A = { insure more than one car } , P ( A ) = 0 . 85. Let B = { insure a sports car } , P ( B ) = 0 . 23. Let C = { insure exactly one car } , P ( C ) = 0 . 15. It is also given that P ( A ∩ B ) = 0 . 17. Since A ∩ C = φ , P ( A ∩ C ) = 0. It follows that P ( A ∩ B ∩ C ′ ) = 0 . 17. Thus P ( A ′ ∩ B ∩ C ′ ) = 0 . 06 and P ( B ′ ∩ C ) = 0 . 09. 1.1-4 (a) S = { HHHH , HHHT , HHTH , HTHH , THHH , HHTT , HTTH , TTHH , HTHT , THTH , THHT , HTTT , THTT , TTHT , TTTH , TTTT } ; (b) (i) 5/16, (ii) 0, (iii) 11/16, (iv) 4/16, (v) 4/16, (vi) 9/16, (vii) 4/16. 1.1-6 (a) P ( A ∪ B ) = 0 . 5 + 0 . 6 − 0 . 4 = 0 . 7; (b) A = ( A ∩ B ′ ) ∪ ( A ∩ B ) P ( A ) = P ( A ∩ B ′ ) + P ( A ∩ B ) 0 . 5 = P ( A ∩ B ′ ) + 0 . 4 P ( A ∩ B ′ ) = 0 . 1; (c) P ( A ′ ∪ B ′ ) = P [( A ∩ B ) ′ ] = 1 − P ( A ∩ B ) = 1 − 0 . 4 = 0 . 6. 1.1-8 Let A = { lab work done } , B = { referral to a specialist } , P ( A ) = 0 . 41 , P ( B ) = 0 . 53 , P ([ A ∪ B ] ′ ) = 0 . 21. P ( A ∪ B ) = P ( A ) + P ( B ) − P ( A ∩ B ) 0 . 79 = 0 . 41 + 0 . 53 − P ( A ∩ B ) P ( A ∩ B ) = 0 . 41 + 0 . 53 − 0 . 79 = 0 . 15 . 1.1-10 A ∪ B ∪ C = A ∪ ( B ∪ C ) P ( A ∪ B ∪ C ) = P ( A ) + P ( B ∪ C ) − P [ A ∩ ( B ∪ C )] = P ( A ) + P ( B ) + P ( C ) − P ( B ∩ C ) − P [( A ∩ B ) ∪ ( A ∩ C )] = P ( A ) + P ( B ) + P ( C ) − P ( B ∩ C ) − P ( A ∩ B ) − P ( A ∩ C ) + P ( A ∩ B ∩ C ) . 1.1-12 (a) 1/3; (b) 2/3; (c) 0; (d) 1/2.
Page 6 Read page 6 as text
2 Section 1.2 Methods of Enumeration 1.1-14 P ( A ) = 2[ r − r ( √ 3 / 2)] 2 r = 1 − √ 3 2 . 1.1-16 Note that the respective probabilities are p 0 , p 1 = p 0 / 4 , p 2 = p 0 / 4 2 , · · · . ∞ ∑ k =0 p 0 4 k = 1 p 0 1 − 1 / 4 = 1 p 0 = 3 4 1 − p 0 − p 1 = 1 − 15 16 = 1 16 . 1.2 Methods of Enumeration 1.2-2 (a) (4)(5)(2) = 40; (b) (2)(2)(2) = 8. 1.2-4 (a) 4 ( 6 3 ) = 80; (b) 4(2 6 ) = 256; (c) (4 − 1 + 3)! (4 − 1)!3! = 20. 1.2-6 S = { DDD, DDFD, DFDD, FDDD, DDFFD, DFDFD, FDDFD, DFFDD, FDFDD, FFDDD, FFF, FFDF, FDFF, DFFF FFDDF, FDFDF, DFFDF, FDDFF, DFDFF, DDFFF } so there are 20 possibilities. Note that the winning player (2 choices) must win the last set and two of the previous sets, so the number of outcomes is 2 [( 2 2 ) + ( 3 2 ) + ( 4 2 )] = 20 . 1.2-8 3 · 3 · 2 12 = 36 , 864 . 1.2-10 ( n − 1 r ) + ( n − 1 r − 1 ) = ( n − 1)! r !( n − 1 − r )! + ( n − 1)! ( r − 1)!( n − r )! = ( n − r )( n − 1)! + r ( n − 1)! r !( n − r )! = n ! r !( n − r )! = ( n r ) . 1.2-12 0 = (1 − 1) n = n ∑ r =0 ( n r ) ( − 1) r (1) n − r = n ∑ r =0 ( − 1) r ( n r ) . 2 n = (1 + 1) n = n ∑ r =0 ( n r ) (1) r (1) n − r = n ∑ r =0 ( n r ) . 1.2-14 ( 5 − 1 + 29 29 ) = 33! 29!4! = 40 , 920. 1.2-16 (a) ( 19 3 )( 52 − 19 6 ) ( 52 9 ) = 102 , 486 351 , 325 = 0 . 2917;
Page 7 Read page 7 as text
Chapter 1 Probability 3 (b) ( 19 3 )( 10 2 )( 7 1 )( 3 0 )( 5 1 )( 2 0 )( 6 2 ) ( 52 9 ) = 7 , 695 1 , 236 , 664 = 0 . 00622 . 1.2-18 (a) P ( A ) = ∑ 5 n =1 (1 / 2) n = 1 − (1 / 2) 5 ; (b) P ( B ) = ∑ 10 n =1 (1 / 2) n = 1 − (1 / 2) 10 ; (c) P ( A ∪ B ) = P ( B ) = 1 − (1 / 2) 10 ; (d) P ( A ∩ B ) = P ( A ) = 1 − (1 / 2) 5 ; (e) P ( C ) = P ( B ) − P ( A ) = (1 / 2) 5 − (1 / 2) 10 ; (f ) P ( B ′ ) = 1 − P ( B ) = (1 / 2) 10 . 1.3 Conditional Probability 1.3-2 (a) 1041 1456 ; (b) 392 633 ; (c) 649 823 . (d) The proportion of women who favor a gun law is greater than the proportion of men who favor a gun law. 1.3-4 (a) P (HH) = 13 52 · 12 51 = 1 17 ; (b) P (HC) = 13 52 · 13 51 = 13 204 ; (c) P (Non-Ace Heart, Ace) + P (Ace of Hearts, Non-Heart Ace) = 12 52 · 4 51 + 1 52 · 3 51 = 51 52 · 51 = 1 52 . 1.3-6 Let H = { died from heart disease } ; P = { at least one parent had heart disease } . P ( H | P ′ ) = N ( H ∩ P ′ ) N ( P ′ ) = 110 648 . 1.3-8 (a) 3 20 · 2 19 · 1 18 = 1 1140 ; (b) ( 3 2 )( 17 1 ) ( 20 3 ) · 1 17 = 1 380 ; (c) 9 ∑ k =1 ( 3 2 )( 17 2 k − 2 ) ( 20 2 k ) · 1 20 − 2 k = 35 76 = 0 . 4605;
Page 8 Read page 8 as text
4 Section 1.4 Independent Events (d) Draw second. The probability of winning is 1 − 0 . 4605 = 0 . 5395. 1.3-10 (a) P ( A ) = 52 52 · 51 52 · 50 52 · 49 52 · 48 52 · 47 52 = 8 , 808 , 975 11 , 881 , 376 = 0 . 74141; (b) P ( A ′ ) = 1 − P ( A ) = 0 . 25859. 1.3-12 (a) It doesn’t matter because P ( B 1 ) = 1 18 , P ( B 5 ) = 1 18 , P ( B 18 ) = 1 18 ; (b) P ( B ) = 2 18 = 1 9 on each draw. 1.3-14 (a) 5 · 4 · 3 = 60; (b) 5 · 5 · 5 = 125. 1.3-16 3 5 · 5 8 + 2 5 · 4 8 = 23 40 . 1.4 Independent Events 1.4-2 (a) P ( A ∩ B ) = P ( A ) P ( B ) = (0 . 3)(0 . 6) = 0 . 18; P ( A ∪ B ) = P ( A ) + P ( B ) − P ( A ∩ B ) = 0 . 3 + 0 . 6 − 0 . 18 = 0 . 72; (b) P ( A | B ) = P ( A ∩ B ) P ( B ) = 0 0 . 6 = 0 . 1.4-4 Proof of (b) : P ( A ′ ∩ B ) = P ( B ) P ( A ′ | B ) = P ( B )[1 − P ( A | B )] = P ( B )[1 − P ( A )] = P ( B ) P ( A ′ ) . Proof of (c) : P ( A ′ ∩ B ′ ) = P [( A ∪ B ) ′ ] = 1 − P ( A ∪ B ) = 1 − P ( A ) − P ( B ) + P ( A ∩ B ) = 1 − P ( A ) − P ( B ) + P ( A ) P ( B ) = [1 − P ( A )][1 − P ( B )] = P ( A ′ ) P ( B ′ ) . 1.4-6 P [ A ∩ ( B ∩ C )] = P [ A ∩ B ∩ C ] = P ( A ) P ( B ) P ( C ) = P ( A ) P ( B ∩ C ) . P [ A ∩ ( B ∪ C )] = P [( A ∩ B ) ∪ ( A ∩ C )] = P ( A ∩ B ) + P ( A ∩ C ) − P ( A ∩ B ∩ C ) = P ( A ) P ( B ) + P ( A ) P ( C ) − P ( A ) P ( B ) P ( C ) = P ( A )[ P ( B ) + P ( C ) − P ( B ∩ C )] = P ( A ) P ( B ∪ C ) . P [ A ′ ∩ ( B ∩ C ′ )] = P ( A ′ ∩ C ′ ∩ B ) = P ( B )[ P ( A ′ ∩ C ′ ) | B ] = P ( B )[1 − P ( A ∪ C | B )] = P ( B )[1 − P ( A ∪ C )] = P ( B ) P [( A ∪ C ) ′ ] = P ( B ) P ( A ′ ∩ C ′ ) = P ( B ) P ( A ′ ) P ( C ′ ) = P ( A ′ ) P ( B ) P ( C ′ ) = P ( A ′ ) P ( B ∩ C ′ ) .
Page 9 Read page 9 as text
Chapter 1 Probability 5 P [ A ′ ∩ B ′ ∩ C ′ ] = P [( A ∪ B ∪ C ) ′ ] = 1 − P ( A ∪ B ∪ C ) = 1 − P ( A ) − P ( B ) − P ( C ) + P ( A ) P ( B ) + P ( A ) P ( C )+ P ( B ) P ( C ) − P A ) P ( B ) P ( C ) = [1 − P ( A )][1 − P ( B )][1 − P ( C )] = P ( A ′ ) P ( B ′ ) P ( C ′ ) . 1.4-8 1 6 · 2 6 · 3 6 + 1 6 · 4 6 · 3 6 + 5 6 · 2 6 · 3 6 = 2 9 . 1.4-10 (a) 3 4 · 3 4 = 9 16 ; (b) 1 4 · 3 4 + 3 4 · 2 4 = 9 16 ; (c) 2 4 · 1 4 + 2 4 · 4 4 = 10 16 . 1.4-12 (a) ( 1 2 ) 3 ( 1 2 ) 2 ; (b) ( 1 2 ) 3 ( 1 2 ) 2 ; (c) ( 1 2 ) 3 ( 1 2 ) 2 ; (d) 5! 3! 2! ( 1 2 ) 3 ( 1 2 ) 2 . 1.4-14 (a) 1 − (0 . 4) 3 = 1 − 0 . 064 = 0 . 936; (b) 1 − (0 . 4) 8 = 1 − 0 . 00065536 = 0 . 99934464. 1.4-16 (a) ∞ ∑ k =0 1 5 ( 4 5 ) 2 k = 5 9 ; (b) 1 5 + 4 5 · 3 4 · 1 3 + 4 5 · 3 4 · 2 3 · 1 2 · 1 1 = 3 5 . 1.4-18 (a) 7; (b) (1 / 2) 7 ; (c) 63; (d) No! (1 / 2) 63 = 1 / 9 , 223 , 372 , 036 , 854 , 775 , 808. 1.4-20 No. The equations that must hold are (1 − p 1 )(1 − p 2 ) = p 1 (1 − p 2 ) + p 2 (1 − p 1 ) = p 1 p 2 . There are no real solutions. 1.5 Bayes’ Theorem 1.5-2 (a) P ( G ) = P ( A ∩ G ) + P ( B ∩ G ) = P ( A ) P ( G | A ) + P ( B ) P ( G | B ) = (0 . 40)(0 . 85) + (0 . 60)(0 . 75) = 0 . 79; (b) P ( A | G ) = P ( A ∩ G ) P ( G ) = (0 . 40)(0 . 85) 0 . 79 = 0 . 43 .
Page 10 Read page 10 as text
6 Section 1.5 Bayes’ Theorem 1.5-4 Let event B denote an accident and let A 1 be the event that age of the driver is 16–25. Then P ( A 1 | B ) = (0 . 1)(0 . 05) (0 . 1)(0 . 05) + (0 . 55)(0 . 02) + (0 . 20)(0 . 03) + (0 . 15)(0 . 04) = 50 50 + 110 + 60 + 60 = 50 280 = 0 . 179 . 1.5-6 Let B be the event that the policyholder dies. Let A 1 , A 2 , A 3 be the events that the deceased is standard, preferred and ultra-preferred, respectively. Then P ( A 1 | B ) = (0 . 60)(0 . 01) (0 . 60)(0 . 01) + (0 . 30)(0 . 008) + (0 . 10)(0 . 007) = 60 60 + 24 + 7 = 60 91 = 0 . 659; P ( A 2 | B ) = 24 91 = 0 . 264; P ( A 3 | B ) = 7 91 = 0 . 077 . 1.5-8 Let A be the event that the tablet is under warranty. P ( B 1 | A ) = (0 . 40)(0 . 10) (0 . 40)(0 . 10) + (0 . 30)(0 . 05) + (0 . 20)(0 . 03) + (0 . 10)(0 . 02) = 40 40 + 15 + 6 + 2 = 40 63 = 0 . 635; P ( B 2 | A ) = 15 63 = 0 . 238; P ( B 3 | A ) = 6 63 = 0 . 095; P ( B 4 | A ) = 2 63 = 0 . 032 . 1.5-10 (a) P ( D + ) = (0 . 02)(0 . 92) + (0 . 98)(0 . 05) = 0 . 0184 + 0 . 0490 = 0 . 0674; (b) P ( A − | D + ) = 0 . 0490 0 . 0674 = 0 . 727; P ( A + | D + ) = 0 . 0184 0 . 0674 = 0 . 273; (c) P ( A − | D − ) = (0 . 98)(0 . 95) (0 . 02)(0 . 08) + (0 . 98)(0 . 95) = 9310 16 + 9310 = 0 . 998; P ( A + | D − ) = 0 . 002; (d) Yes, particularly those in part (b). 1.5-12 Let D = { defective roll } . Then P ( I | D ) = P ( I ∩ D ) P ( D ) = P ( I ) · P ( D | I ) P ( I ) · P ( D | I ) + P ( II ) · P ( D | II ) = (0 . 60)(0 . 03) (0 . 60)(0 . 03) + (0 . 40)(0 . 01) = 0 . 018 0 . 018 + 0 . 004 = 0 . 018 0 . 022 = 0 . 818 .
Page 11 Read page 11 as text
Chapter 2 Discrete Distributions 7 Chapter 2 Discrete Distributions 2.1 Random Variables of the Discrete Type 2.1-2 (a) f ( x ) = 0 . 6 , x = 1 , 0 . 3 , x = 5 , 0 . 1 , x = 10; (b) f(x) x 0.1 0.2 0.3 0.4 0.5 0.6 1 2 3 4 5 6 7 8 9 10 Figure 2.1–2: A probability histogram 2.1-4 (a) 9 ∑ x =1 log 10 ( x + 1 x ) = 9 ∑ x =1 [log 10 ( x + 1) − log 10 x ] = log 10 2 − log 10 1 + log 10 3 − log 10 2 + · · · + log 10 10 − log 10 9 = log 10 10 = 1; (b) F ( x ) = 0 , x < 1 , log 10 n, n − 1 ≤ x < n, n = 2 , 3 , . . . , 9 , 1 , 9 ≤ x.
Page 12 Read page 12 as text
8 Section 2.1 Random Variables of the Discrete Type 2.1-6 (a) f ( x ) = 1 10 , x = 0 , 1 , 2 , · · · , 9; (b) N ( { 0 } ) / 150 = 11 / 150 = 0 . 073; N ( { 5 } ) / 150 = 13 / 150 = 0 . 087; N ( { 1 } ) / 150 = 14 / 150 = 0 . 093; N ( { 6 } ) / 150 = 22 / 150 = 0 . 147; N ( { 2 } ) / 150 = 13 / 150 = 0 . 087; N ( { 7 } ) / 150 = 16 / 150 = 0 . 107; N ( { 3 } ) / 150 = 12 / 150 = 0 . 080; N ( { 8 } ) / 150 = 18 / 150 = 0 . 120; N ( { 4 } ) / 150 = 16 / 150 = 0 . 107; N ( { 9 } ) / 150 = 15 / 150 = 0 . 100. (c) x f(x), h(x) 0.02 0.04 0.06 0.08 0.10 0.12 0.14 1 2 3 4 5 6 7 8 9 Figure 2.1–6: Michigan daily lottery digits 2.1-8 (a) f ( x ) = 6 − | 7 − x | 36 , x = 2 , 3 , 4 , 5 , 6 , 7 , 8 , 9 , 10 , 11 , 12; (b) x f(x) 0.02 0.04 0.06 0.08 0.10 0.12 0.14 0.16 1 2 3 4 5 6 7 8 9 10 11 12 Figure 2.1–8: Probability histogram for the sum of a pair of dice
Page 13 Read page 13 as text
Chapter 2 Discrete Distributions 9 2.1-10 (a) The space of W is S = { 0 , 1 , 2 , 3 , 4 , 5 , 6 , 7 } . P ( W = 0) = P ( X = 0 , Y = 0) = 1 2 · 1 4 = 1 8 , assuming independence. P ( W = 1) = P ( X = 0 , Y = 1) = 1 2 · 1 4 = 1 8 , P ( W = 2) = P ( X = 2 , Y = 0) = 1 2 · 1 4 = 1 8 , P ( W = 3) = P ( X = 2 , Y = 1) = 1 2 · 1 4 = 1 8 , P ( W = 4) = P ( X = 0 , Y = 4) = 1 2 · 1 4 = 1 8 , P ( W = 5) = P ( X = 0 , Y = 5) = 1 2 · 1 4 = 1 8 , P ( W = 6) = P ( X = 2 , Y = 4) = 1 2 · 1 4 = 1 8 , P ( W = 7) = P ( X = 2 , Y = 5) = 1 2 · 1 4 = 1 8 . That is, f ( w ) = P ( W = w ) = 1 8 , w ∈ S . (b) w ( ) f w 0.02 0.04 0.06 0.08 0.10 0.12 1 2 3 4 5 6 7 Figure 2.1–10: Probability histogram of sum of two special dice 2.1-12 Let x equal the number of orange balls and 144 − x the number of blue balls. Then x 144 · x − 1 143 + 144 − x 144 · 143 − x 143 = x 144 · 144 − x 143 + 144 − x 144 · x 143 x 2 − x + 144 · 143 − 144 x − 143 x + x 2 = 2 · 144 x − 2 · x 2 x 2 − 144 x + 5 , 148 = 0 ( x − 78)( x − 66) = 0 Thus there are 78 orange balls and 66 blue balls. 2.2 Mathematical Expectation 2.2-2 E ( X ) = ( − 1) ( 4 9 ) + (0) ( 1 9 ) + (1) ( 4 9 ) = 0; E ( X 2 ) = ( − 1) 2 ( 4 9 ) + (0) 2 ( 1 9 ) + (1) 2 ( 4 9 ) = 8 9 ;
Page 14 Read page 14 as text
10 Section 2.3 Special Mathematical Expectations E (3 X 2 − 2 X + 4) = 3 ( 8 9 ) − 2(0) + 4 = 20 3 . 2.2-4 1 = 6 ∑ x =0 f ( x ) = 9 10 + c ( 1 1 + 1 2 + 1 3 + 1 4 + 1 5 + 1 6 ) c = 2 49 ; E (Payment) = 2 49 ( 1 · 1 2 + 2 · 1 3 + 3 · 1 4 + 4 · 1 5 + 5 · 1 6 ) = 71 490 units . 2.2-6 Note that ∞ ∑ x =1 6 π 2 x 2 = 6 π 2 ∞ ∑ x =1 1 x 2 = 6 π 2 π 2 6 = 1, so this is a pdf. E ( X ) = ∞ ∑ x =1 x 6 π 2 x 2 = 6 π 2 ∞ ∑ x =1 1 x = + ∞ and it is well known that the sum of this harmonic series is not finite. 2.2-8 E ( | X − c | ) = 1 7 ∑ x ∈ S | x − c | , where S = { 1 , 2 , 3 , 5 , 15 , 25 , 50 } . When c = 5, E ( | X − 5 | ) = 1 7 [(5 − 1) + (5 − 2) + (5 − 3) + (5 − 5) + (15 − 5) + (25 − 5) + (50 − 5)] . If c is either increased or decreased by 1, this expectation is increased by 1 / 7. Thus c = 5, the median, minimizes this expectation while b = E ( X ) = 101 / 7 minimizes E [( X − b ) 2 ]. 2.2-10 (1) · 15 36 + ( − 1) · 21 36 = − 6 36 = − 1 6 ; (1) · 15 36 + ( − 1) · 21 36 = − 6 36 = − 1 6 ; (4) · 6 36 + ( − 1) · 30 36 = − 6 36 = − 1 6 . 2.2-12 (a) The average class size is (16)(25) + (3)(100) + (1)(300) 20 = 50; (b) f ( x ) = 0 . 4 , x = 25 , 0 . 3 , x = 100 , 0 . 3 , x = 300; (c) E ( X ) = 25(0 . 4) + 100(0 . 3) + 300(0 . 3) = 130 .
Page 15 Read page 15 as text
Chapter 2 Discrete Distributions 11 2.3 Special Mathematical Expectations 2.3-2 (a) μ = E ( X ) = 3 ∑ x =1 x 3! x ! (3 − x )! ( 1 4 ) x ( 3 4 ) 3 − x = 3 ( 1 4 ) 2 ∑ k =0 2! k ! (2 − k )! ( 1 4 ) k ( 3 4 ) 2 − k = 3 ( 1 4 )( 1 4 + 3 4 ) 2 = 3 4 ; E [ X ( X − 1)] = 3 ∑ x =2 x ( x − 1) 3! x ! (3 − x )! ( 1 4 ) x ( 3 4 ) 3 − x = 2(3) ( 1 4 ) 2 3 4 + 6 ( 1 4 ) 3 = 6 ( 1 4 ) 2 = 2 ( 1 4 )( 3 4 ) ; σ 2 = E [ X ( X − 1)] + E ( X ) − μ 2 = (2) ( 3 4 )( 1 4 ) + ( 3 4 ) − ( 3 4 ) 2 = (2) ( 3 4 )( 1 4 ) + ( 3 4 )( 1 4 ) = 3 ( 1 4 )( 3 4 ) ; (b) μ = E ( X ) = 4 ∑ x =1 x 4! x ! (4 − x )! ( 1 2 ) x ( 1 2 ) 4 − x = 4 ( 1 2 ) 3 ∑ k =0 3! k ! (3 − k )! ( 1 2 ) k ( 1 2 ) 3 − k = 4 ( 1 2 )( 1 2 + 1 2 ) 3 = 2; E [ X ( X − 1)] = 4 ∑ x =2 x ( x − 1) 4! x ! (4 − x )! ( 1 2 ) x ( 1 2 ) 4 − x = 2(6) ( 1 2 ) 4 + (6)(4) ( 1 2 ) 4 + (12) ( 1 2 ) 4 = 48 ( 1 2 ) 4 = 12 ( 1 2 ) 2 ; σ 2 = (12) ( 1 2 ) 2 + 4 2 − ( 4 2 ) 2 = 1 . 2.3-4 E [( X − μ ) /σ ] = (1 /σ )[ E ( X ) − μ ] = (1 /σ )( μ − μ ) = 0; E { [( X − μ ) /σ ] 2 } = (1 /σ 2 ) E [( X − μ ) 2 ] = (1 /σ 2 )( σ 2 ) = 1. 2.3-6 f (1) = 3 8 , f (2) = 2 8 , f (3) = 3 8 ,
Page 16 Read page 16 as text
12 Section 2.3 Special Mathematical Expectations μ = 1 · 3 8 + 2 · 2 8 + 3 · 3 8 = 2, σ 2 = 1 2 · 3 8 + 2 2 · 2 8 + 3 2 · 3 8 − 2 2 = 3 4 . 2.3-8 E ( X ) = 4 ∑ x =1 x · 2 x − 1 16 = 50 16 = 3 . 125; E ( X 2 ) = 4 ∑ x =1 x 2 · 2 x − 1 16 = 85 8 ; Var( X ) = 85 8 − ( 25 8 ) 2 = 55 64 = 0 . 8594; σ = √ 55 8 = 0 . 9270 . 2.3-10 μ = E ( X ) = ( − 5)(1 / 16) + ( − 1)(5 / 8) + (3)(5 / 16) = 0 so E [( X − μ ) 3 ] = E ( X 3 ) = ( − 5) 3 (1 / 16)+( − 1) 3 (5 / 8)+(3 3 )(5 / 16) = ( − 125 − 10+135) / 16 = 0. Thus the index of skewness likewise equals 0. The distribution is not symmetric, however. 2.3-12 (a) f ( x ) = ( 364 365 ) x − 1 ( 1 365 ) , x = 1 , 2 , 3 , . . . ; (b) μ = 1 1 365 = 365 , σ 2 = 364 365 ( 1 365 ) 2 = 132 , 860 , σ = 364 . 500; (c) P ( X > 400) = ( 364 365 ) 400 = 0 . 3337 , P ( X < 300) = 1 − ( 364 365 ) 299 = 0 . 5597 . 2.3-14 P ( X ≥ 100) = P ( X > 99) = (0 . 99) 99 = 0 . 3697 . 2.3-16 (a) f ( x ) = (1 / 2) x − 1 , x = 2 , 3 , 4 , . . . ; (b) M ( t ) = E [ e tx ] = ∞ ∑ x =2 e tx (1 / 2) x − 1 = 2 ∞ ∑ x =2 ( e t / 2) x = 2( e t / 2) 2 1 − e t / 2 = e 2 t 2 − e t , t < ln 2;
127 more pages in the full document
Unlock it once and it stays in your library, ready to chat with or turn into flashcards.
Unlock the full documentMore Statistics study materials
See all
Exam GuidesTest Bank for Biostatistics for the Biological and Health Sciences , 2nd Edition
Statistics
182 pages161downloadsFeb 2025
Exam GuidesElementary Statistics, 13th Edition Test Bank
Statistics
239 pages160downloadsFeb 2025
Textbook GuidesSolution Manual For Statistics for Managers Using Microsoft Excel, 7th Edition
Statistics
1418 pages158downloadsFeb 2025
Textbook GuidesStatistics for Managers Using Microsoft Excel 8th Edition Solution Manual
Statistics
1343 pages156downloadsFeb 2025
Textbook GuidesEssentials of Social Statistics for a Diverse Society Second Edition Solution Manual
Statistics
137 pages135downloadsFeb 2025
Exam GuidesTest Bank For Elementary Statistics, 3rd Edition
Statistics
969 pages126downloadsFeb 2025
Textbook GuidesSolution Manual for Business Statistics, 4th Edition
Statistics
582 pages125downloadsFeb 2025
Exam GuidesTest Bank for Business Statistics: A Decision-Making Approach, 10th Edition
Statistics
881 pages110downloadsFeb 2025
Study this document with CramX
Once it is in your library, every tool can work from it.

